### v2

# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

Abstract
In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor. This study expands the temporal difference decomposition approach, TD($\Delta$), to the SARSA algorithm, which we designate as SARSA($\Delta$). SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates. TD($\Delta$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors. This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization. We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in both deterministic and stochastic environments. Experimental findings across many benchmark tasks indicate that the proposed SARSA($\Delta$) surpasses conventional TD learning methods in both tabular and deep RL environments.
keywords:
SARSA, Temporal Difference, Action-Value Functions, On-policy, Finite Horizon
1 Introduction
In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment. Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards. Traditionally, these strategies utilise a discount factor $10\leq\eta<$ 1, with values approaching $\eta=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1]. Discount factors $\eta<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2]. In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent. The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\eta$ employed in SARSA results in a compromise between bias and variance. Small values of $\eta$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient. In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
Recent research in TD learning has investigated the segmentation of value functions across multiple time scales, enabling agents to enhance value estimates concurrently at different horizons[8]. The TD($\Delta$) architecture improves scalability and performance in scenarios involving long-term dependencies by breaking down value functions into components that incorporate different discount factors, known as TD($\Delta$). The adaptability of the TD($\Delta$) framework shows considerable promise in addressing the bias-variance trade-offs that are characteristic of TD learning.
We propose SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle. TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators. These estimators approximate the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions, thereby improving learning efficiency in scenarios where actions affect long-horizon outcomes. Similar to TD($\Delta$), SARSA($\Delta$) seeks to optimize a set of delta estimators, where each estimator is associated with a distinct discount factor. With smaller discount factors, convergence happens faster because to segmentation, while greater discount factors strengthen this foundation for better long-horizon planning.
This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates action-value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate deterministic and stochastic environments. Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces. Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation. SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
The impetus for extending TD($\Delta$) to SARSA($\Delta$) lies in the fact that temporal difference (TD($\Delta$)) learning offers a framework for deconstructing value functions across many time scales, hence overcoming the constraints of single-scale learning in long-horizon problems. By extending TD($\Delta$) to SARSA($\Delta$), we seek to decompose the action-value function Q(s,a) into elements that emphasise distinct discount factors. This decomposition enables SARSA to utilise the advantages of learning across several time scales, hence stabilising the learning process and mitigating the bias-variance tradeoff. This addon provides a more detailed and scalable method for action-value function learning in intricate environments.
The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time-scales. (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function. (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
The ensuing portions of the paper are organized as follows: Section 2 presents a discourse on pertinent literature. Section 3 provides the essential background information. The principles of SARSA($\Delta$) and theoretical analysis are shown in Sections 4 and 5, respectively. Section 6 provides a detailed summary of the experiments performed. Ultimately, we present a conclusion in section 7.
2 Related Work
In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return. Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias. Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance. The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in RL. The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL. Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors. The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods. Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based RL algorithms such as Rainbow. This method demonstrates potential in contexts characterized by uncertainty and risk. Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning. This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result. The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial. This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation. Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule. It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions. Our method can also be utilised as a generalised action-value function, and its capacity to query smaller time-scales represents an additional advantage. Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
Recently, there has been some research conducted on the precise selection of the discount factor[2, 5, 6, 4]. The authors[6] of this study introduce meta-gradient methods for the adaptive selection of discount factors in RL, emphasising the necessity of balancing short- and long-term rewards, which corresponds with the rationale for dissecting action-value functions across several discount factors in SARSA($\Delta$).
Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance trade-off. These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales. This substantiates the assertion that value function decomposition can enhance learning efficacy and stability. This paper[9] functions as a theoretical foundation for comprehending the bias-variance trade-off that arises when decomposing Q-values across various time-scales, and it also establishes constraints on the bias and variance of TD updates. This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments. It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $n$, the agent occupies state $s_{n}$, chooses an action $a_{n}$, obtains a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the subsequent state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent. The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\eta}$ can approximate the true action-value function $Q_{\eta}$. The one-step TD error $\delta^{\eta}_{n}=r_{n+1}+\eta\hat{Q}_{\eta}(s_{n+1},a_{n+1})\mathit{-}\hat{Q}_{\eta}(s_{n},a_{n})$ is used to update the action-value function given a transition $(s_{n},a_{n},r_{n},s_{n+1})$.
An on-policy RL technique called SARSA learns the action-value function Q(s,a) for a given policy $\pi$. The action-value function represents the expected total reward that comes from state s, taking action a, and following policy $\pi$. Then the SARSA update rule is clarified as follows:
$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (1)
Here, $\alpha$ is the learning rate, $\eta$ is the discount factor, and $r_{n}$ is the immediate reward. Long-horizon tasks present challenges for SARSA since the choice of $\eta$ can compromise learning efficiency with appropriate long-term reward maximizing.
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24]. The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\eta$. The equation for the Generalized Advantage Estimator is generally expressed as:
$A(s_{n},a_{n})=\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{t+k}^{\eta}$ (2)
Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$
Loss function for Q-Value estimation utilizing GAE. The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{n},a_{n};\theta)$ and the target Q-value adjusted by the advantage estimator. Thus, it is possible to concisely write the loss function using $A(s_{n},a_{n})$ and $Q(s_{n},a_{n})$ as follows:
$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{n},a_{n};\theta)-\bigg{(}Q(s_{n},a_{n})+A(s_{n},a_{n})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4. For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\nu}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}-log\pi(a|s;\nu)A(s,a)\bigg{]}$ (4)
Proximal Policy Optimization (PPO) [28], an enhancement of actor-critic methodologies, confines policy updates to a designated optimization area, referred to as a trust region, utilizing a clipping objective that evaluates the current parameters, $\nu$, against the prior parameters, $\nu_{old}$:
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),\psi(\nu)A(s,a)\bigg{)}\bigg{]}$ (5)
In on-policy methods like SARSA, the likelihood ratio $\rho(\nu)$ between the current policy and a prior policy is typically employed. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. However, when utilized within an actor-critic framework, the SARSA-learning agent may adopt a similar policy ratio:
$\rho(\nu)=\frac{Q(s,a;\nu)}{Q(s,a;\nu_{old})}$ (6)
where $\psi(\nu)=clip(\rho,1-\epsilon,1+\epsilon)$ represents the clipped likelihood ratio, and $\epsilon<1$ is a negligible parameter employed to constrain the update. The loss function for the policy update can be articulated as follows:
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A(s,a)\bigg{)}\bigg{]}$ (7)
4 Methodology
The TD($\Delta$) paradigm boosts regular TD learning by splitting action-value functions over several discount factors $\eta_{0},\eta_{1},...,\eta_{m}$. This decomposition facilitates the derivation of the action-value function as a sum of delta estimators, where each estimator indicates the distinction between action-value functions at various discount factors. Concentrating on smaller, more manageable time scales facilitates the regulation of learning variance and bias, which constitutes the primary benefit of this approach.
4.1 TD($\Delta$) Framework
To compute $D_{m}$ for SARSA as outlined in the study, it is essential to comprehend the mechanism by which delta estimators ($D_{m}$) approximate the discrepancies between action-value functions across successive discount factors. This approach for SARSA will be outlined in a stepwise manner.
Delta Estimators ($D_{m}$) and Action-Value Functions ($Q(s,a)$): Delta estimators, denoted as $D_{m}$, quantify the variation between action-value functions associated with consecutive discount factors:
$D_{m}=Q_{\eta_{m}}\mathit{-}Q_{\eta_{m-1}}$ (8)
In this context, $Q_{\eta_{m}}$ denotes the action-value function with discount factor $\eta_{m}$, but $Q_{\eta_{m-1}}$ signifies the action-value function with the prior discount factor, $\eta_{m-1}$.
4.2 Single-Step TD ( SARSA($\Delta$) )
This section initiates with an overview of the delta estimators ($D_{m}$) pertaining to the action-value functions $Q$ employed in SARSA. SARSA is extended to TD($\Delta$), which is referred to as SARSA($\Delta$). SARSA($\Delta$) seeks to derive the action-value function $Q(s,a)$, which indicates the anticipated return of following policy $\pi$ from state $s$, executing action $a$, and then persistently adhering to $\pi$ thereafter. The objective is to enhance the total reward by optimizing the action-value function over various time-scales. The fundamental concept in SARSA($\Delta$) is the segmentation of the action-value function $Q(s,a)$ into multiple delta components $D_{m}(s,a)$, where each component is linked to a distinct discount factor $\eta_{m}$. The relationship between these delta components (i.e., delta function) is expressed in the following manner:
$D_{m}(s,a):=Q_{\eta_{m}}(s,a)\mathit{-}Q_{\eta_{m-1}}(s,a)$ (9)
where $\eta_{0},\eta_{1},...,\eta_{m}$ represent the discount factors across various time-scales, and define $D_{0}(s,a):=Q_{\eta_{0}}(s,a)$. The action-value function $Q_{\eta_{m}}(s,a)$ is defined as the cumulative sum of all D-components up to m:
$Q_{\eta_{m}}(s,a)=\sum_{x=0}^{m}D_{x}(s,a)$ (10)
The update rule in standard SARSA is determined by the state-action pair in accordance with the policy. The action-value function is updated through the TD error, with the action in the subsequent state selected in accordance with the policy:
$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (11)
In this context, $r_{n}$ signifies the reward acquired from performing action $a_{n}$ in state $s_{n}$, whereas $(s_{n+1},a_{n+1})$ denotes the ensuing state-action pair. The SARSA update rule is implemented independently to each delta component $D_{n}(s,a)$. The modifications for each component utilize the same architecture as the conventional SARSA approach, but it is tailored to include the delta function at each distinct time-scale. The update in Eq.11 for single-step TD SARSA ($\Delta$) can be articulated with numerous time-scales by employing the principle of splitting the update into several discount factors[8] as demonstrated below:
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (12)
where $Q_{\eta_{m}}(s_{n},a_{n})$ represents the action-value function with a discount factor $\eta_{m}$. $D_{m}(s_{n},a_{n})$ denotes the delta function for SARSA ($\Delta$), while $\eta_{m}$ represents the discount factor for time-scale m, while $Q_{\eta_{m}}(s_{n+1},a_{n+1})$ is the SARSA update for action-value associated with the subsequent state-action pair. The single-step TD SARSA ($\Delta$) mirrors the original TD update, but it is applied to action-value functions Q rather than state-value functions V. The single-step Bellman Eq. for the Q-value function in standard SARSA is as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[r_{n}+\eta_{n}Q_{\eta_{n}}(s_{n+1},a_{n+1})\right]$ (13)
This denotes the expected value of the reward acquired at time n, in addition to the discounted value of the action-value function in the subsequent state-action pair $(s_{n+1},a_{n+1})$. Currently, rather than utilizing a singular discount factor $\eta$, we are employing a sequence of discount factors $\eta_{0},\eta_{1},.....,\eta_{m}$. The delta function $D_{m}(s_{n},a_{n})$ is defined as the difference between action-value functions associated with successive discount factors from Eq. 9:
$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$ (14)
To expand the delta function $D_{z}$ utilizing its definition, we replace the Bellman Eq. for both $Q_{\eta_{m}}$ and $Q_{\eta_{m-1}}$, thereby expressing $D_{m}$ in relation to these action-value functions. Commence by composing:
[…]
Since the reward terms $r_{n}$ are included in both formulations, simplify by eliminating them.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}Q_{\eta_{m}}(s_{n+1},a_{n+1})-\eta_{m-1}Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
Decomposing the terms allows for a more detailed analysis of the expression. Utilizing the recursive relationship from Eq. 14
$Q_{\eta_{m}}(s_{n+1},a_{n+1})=D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$
Substitute the definition of $D_{m}$ into the Eq. below.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}(D_{m}(s_{n+1},a_{n+1})+Q_{\eta_{m-1}}(s_{n+1},a_{n+1}))-\eta_{n-1}Q_{\eta_{n-1}}(s_{n+1},a_{n+1})\right]$
Make the terms simpler:
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\eta_{m}D_{m}(s_{n+1},a_{n+1})+(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})\right]$
Integrate the terms to derive the final update Eq. for $D_{m}(s_{n},a_{n})$.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[(\eta_{m}-\eta_{m-1})Q_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}D_{m}(s_{n+1},a_{n+1})\right]$ (16)
This Eq. demonstrates that the update for $D_{m}$ integrates the disparity in discount factors $\eta_{m}-\eta_{m-1}$ with respect to the action-value function $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ and the bootstrapping from the subsequent step $D_{m}(s_{n+1},a_{n+1})$. This denotes a Bellman Eq. for $D_{m}$, combining a decay factor $\eta_{m}$ and the reward $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$. Consequently, it can be utilized to define the expected TD update for $D_{m}$. In this expression, $Q_{\eta_{m-1}}(s_{n+1},a_{n+1})$ can be expressed as the summation of $D_{m}(s_{n+1},a_{n+1})$ for $m\leq M-1$, indicating that the Bellman Eq. for $D_{m}$ is contingent upon the values of all delta functions $D_{m}$ for $m\leq M-1$.
This approach treats the delta value function at each time-scale as an autonomous RL issue, with rewards obtained from the action-value function of the immediately preceding time-scale. Thus, for a target discounted action-value function $Q_{\eta_{m}}(s,a)$, all delta components can be trained simultaneously through a TD update, utilizing the prior values of each estimator for bootstrapping. This process necessitates the assumption of a sequence of discount factors, denoted as $\eta_{m}$, which includes both the minimum and maximum values, $\eta_{0}$ and $\eta_{m}$[8].
4.3 Multi-Step TD ( SARSA($\Delta$) )
Numerous studies indicate that multi-step TD methods typically exhibit greater efficiency compared to single-step TD methods [29]. In multi-step TD SARSA($\Delta$), rewards are aggregated over multiple steps instead of depending on a single future reward. This approach takes into account the differences between consecutive discount factors, $\eta_{m}$ and $\eta_{m-1}$, while utilizing the value function at state $s_{n+k_{m}}$ for bootstrapping. In SARSA($\Delta$), the update Eqs. are modified to aggregate rewards across a series of transitions prior to bootstrapping from either the Q-values or delta estimators. In multi-step SARSA($\Delta$), the agent aggregates rewards over K steps, leading to a corresponding modification in the update rule for each $D_{m}$.
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=1}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{x}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (17)
In this case, $r_{m+x}$ represents the reward obtained at time step $n+x$. $a_{n+k_{m}}$ and the action is chosen greedily as $a=\arg\max_{a}Q(s_{n+k_{m}},a)$, where $k_{m}$ is the number of steps linked to the discount factor $\eta_{m}$. This method generalizes Q-learning to multi-step TD learning by aggregating rewards across several stages and utilizing bootstrapping from both the current and prior time-scale action-value functions. In normal multi-step TD learning, the TD error is computed by aggregating rewards over several stages rather than relying on a single future step, followed by bootstrapping from the action-value at the last step. For example, utilizing a singular discount factor $\eta$, the multi-step TD update equation is expressed as:
$Q(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k-1}\eta^{x}r_{n+x}+\eta^{k}Q(s_{n+k},a_{n+k})\right]$ (18)
Herein, The initial segment of the Eq. aggregates the rewards across k steps, each diminished by $\eta$. The second component derives from the action-value function at time step n+k. We expand this methodology to include various discount factors and delta decomposition, concentrating on representing the delta function $D_{m}(s_{n},a_{n})$ as the difference between action-value functions with discount factors $\eta_{m}$ and $\eta_{m-1}$. According to the definition of $D_{m}$, we obtain:
$D_{m}(s_{n},a_{n})=Q_{\eta_{m}}(s_{n},a_{n})-Q_{\eta_{m-1}}(s_{n},a_{n})$
We now implement the multi-step variant of the Bellman Eq. for both $Q_{\eta_{m}}(s_{n},a_{n})$ and $Q_{\eta_{m-1}}(s_{n},a_{n})$. The multi-step Bellman Eq. for $Q_{\eta_{m}}(s_{n},a_{n})$ is expressed as follows:
$Q_{\eta_{m}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
For $Q_{\eta_{m-1}}(s_{n},a_{n})$, the multi-step Bellman Eq. is expressed as follows:
$Q_{\eta_{m-1}}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
We will now utilize the delta estimator formulas to subtract these two Eqs.
$D_{m}(s_{n},a_{n})=\left[\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}+\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})\right]-\left[\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}+\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})\right]$
We will now broaden and simplify the expressions. The immediate reward terms $r_{n+x}$ are present in both Eqs. but are adjusted by distinct discount factors, $\eta_{m}$ and $\eta_{m-1}$. This enables us to express the disparity in rewards as:
$\sum_{x=0}^{k_{m}-1}\eta_{m}^{x}r_{n+x}-\sum_{x=0}^{k_{m}-1}\eta_{m-1}^{x}r_{n+x}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}$
Subsequently, for the bootstrapping terms, we utilize the recursive relationship as outlined:
$Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})=D_{m}(s_{n+k_{m}},a_{n+k_{m}})+Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$
As a result, we get:
$\eta_{m}^{k_{m}}Q_{\eta_{m}}(s_{n+k_{m}},a_{n+k_{m}})-\eta_{m-1}^{k_{m}}Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})=(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
We now amalgamate the terms for both the rewards and bootstrapping to produce the final expression:
$D_{m}(s_{n},a_{n})=\mathbb{E}\left[\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})\right]$ (19)
This Eq. extends the conventional multi-step TD update to include various multiple discount factors and action-value functions in RL. As a result, each $D_{m}$ obtains a share of the rewards from the environment up to time-step $k_{m}-1$. Furthermore, each $D_{m}$ utilizes its distinct action-value function in conjunction with the value from the prior time-scale. A variation of this algorithm, utilizing k-step bootstrapping as described in [29], is presented in Algorithm 1. Despite Algorithm 1 exhibiting quadratic complexity in relation to M, it can be executed with linear complexity for substantial M by preserving $\hat{Q}$ action-values at each time-scale $\eta_{m}$.
Algorithm 1 Multi-Step TD ( SARSA($\Delta$) )
Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
Set the initial value of $D_{m}(s,a)=0$ for all states, actions, and scales m.
for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
     Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
         Select action $a_{n+1}$ based on a policy.
         for m = 0, 1, …, M do
              if m = 0 then
                  $G^{0}=\sum_{x=0}^{k_{0}-1}\eta_{0}^{x}r_{n+x}+\eta_{0}^{k_{0}}D_{0}(s_{n+k_{0}},a_{n+k_{0}})$
              $\triangleright$ Utilizing Eq.10, we substitute $Q_{\eta_{m-1}}(s_{n+k_{m}},a_{n+k_{m}})$ by summing the D-components up to $D_{m-1}$ in Eq.4.3.
                  $G^{m}=\sum_{x=0}^{k_{m}-1}(\eta_{m}^{x}-\eta_{m-1}^{x})r_{n+x}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\sum_{m=0}^{m-1}D_{m-1}(s_{n+k_{m}},a_{n+k_{m}})+\eta_{m}^{k_{m}}D_{m}(s_{n+k_{m}},a_{n+k_{m}})$
              end if
         end for
         for m = 0, 1, 2,…, M do
              $D_{m}(s_{n},a_{n})\leftarrow D_{m}(s_{n},a_{n})+\alpha_{m}(G^{m}-D_{m}(s_{n},a_{n}))$
         end for
     end for
end for
4.4 SARSA TD($\lambda,\Delta$)
Eq.20 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is defined as:
$G_{n}^{\eta,\lambda}(s_{n},a_{n})=\hat{Q}_{\eta}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{n+k}^{\eta}$ (20)
Eq.21 defines the TD($\lambda$) operator, which is used to iteratively apply $\lambda$-returns in updating the value functions. For SARSA, the TD($\lambda$) operator updates the action-value function by summing the $\lambda$-discounted TD errors as follows:
$\mathrm{T}_{\lambda}Q(s_{n},a_{n})=Q(s_{n},a_{n})+(I-\lambda\eta P)^{-1}(\mathrm{T}Q(s_{n},a_{n})-Q(s_{n},a_{n}))$ (21)
Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function. Similarly, Eq.22 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
$G_{n}^{m,\lambda_{m}}:=\hat{D}_{m}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{m}$ (22)
where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})-\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, we compute advantage estimates $A^{\Delta}(s_{n},a_{n})$ for each time-scale. Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
$A^{\Delta}(s_{n},a_{n})=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{\Delta}$ (23)
where $\delta_{n+k}^{\Delta}:=r_{n}+\eta_{m}\sum_{m=0}^{M}\hat{D}_{m}(s_{n+1},a_{n+1})-\sum_{m=0}^{M}\hat{D}_{m}(s_{n},a_{n})$.
We utilize $\eta_{m}$ as the discount factor and the sum of all D estimators as a surrogate for $Q_{\eta m}$. This objective is applicable to PPO by implementing the policy update from Eq.7 and replacing A with $A^{\Delta}$. Additionally, to train each $D_{m}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.22. For more information, see Algorithm 2.
Algorithm 2 PPO-TD($\lambda$, SARSA($\Delta$))
Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
Set the initial value of $D_{m}(s,a)=0$ for all states s, actions a, and scales m.
Initialize policy $\nu$, and values $\theta^{m}\forall m$
for episode = 0, 1, 2,… $\triangleright$ Loop for each episode
     Initialize state $s_{0}$ and choose an initial action $a_{0}$ in accordance with a policy.
     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
         Select next action $a_{n+1}$ based on a policy.
         for m = 0, 1, …, M do
              if $n\geq T$ then
                  $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$triangleright Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.22.
              end if
         end for
         for m = 0, 1, 2,…, M do
              $\hat{D}_{m}(s_{n-T},a_{n-T})\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)$
              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq.23.
              $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$triangleright Update $\theta^{m}$ with TD (Eq.3) using $G^{m,\lambda_{m}}\forall m$.
              $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$triangleright from Eq.7
              $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$triangleright Update the policy parameters $\nu$ with PPO (Eq.7) for SARSA using $A^{\Delta}$
         end for
     end for
end for
5 Analysis
Next, We examine our delta estimators through the lens of the bias-variance trade-off. In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\eta$ in $D_{m}$. Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards. In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability. In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\eta}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator. Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results[8].
5.1 SARSA Equivalency Configurations and Enhancements
In specific situations, we can show that our delta estimators for SARSA align with the conventional action-value estimator when restructured into an action-value function. This discussion centers on linear function approximation of the specified form:
$\hat{Q}(s,a)_{\eta}:=\langle\theta^{\eta},\phi(s,a)\rangle\quad\text{and}\quad\hat{D}_{m}(s,a):=\langle\theta^{m},\phi(s,a)\rangle,\forall m$
where $\theta$ and $\theta^{m}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space. The weight vector $\theta$ for SARSA is updated according to the TD($\lambda$) learning rule in the following manner:
$\theta_{n+1}^{\eta}=\theta_{n}^{\eta}+\alpha\bigg{(}G_{n}^{\eta,\lambda}-\hat{Q}_{\eta}(s,a)\bigg{)}\phi(s_{n},a_{n}),$ (24)
where $G_{n}^{\eta,\lambda}$ represents the TD($\lambda$) return defined in Eq.20. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
$\theta_{n+1}^{m}=\theta_{n}^{m}+\alpha\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n}),$ (25)
where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.22, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates. The following theorem demonstrates the equivalence of two SARSA version algorithms in this context. The subsequent theorems resemble those presented by Romoff et al. [8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
Theorem 1.
Inspired by Romoff et al.[8], if $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$, $\forall m$, and if we choose the initial conditions in such a way that $\sum_{m=0}^{M}\theta_{0}^{\eta}=\theta_{0}^{\eta}$, then the iterates produced by TD($\lambda$) (Eq. 24) and TD($\lambda,\Delta$) (Eq. 25) with linear function approximation satisfy the following conditions:
$\Sigma_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$ (26)
The proof is presented in appendix PROOFS.
Equivalence is attained when $\lambda_{m}\eta_{m}=\lambda_{\eta},\forall m$. When $\lambda$ approaches 1 and $\eta_{m}$ is less than $\eta$, this condition suggests that $\lambda_{m}=\lambda_{\eta}/\eta_{m}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{m}$). The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.21, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\eta}{2\eta}$, thereby confirming that $\lambda_{\eta}<1$[8].
Theorem 2.
Inspired by Romoff et al.[8], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
The proof is presented in appendix PROOFS.
The analysis presented in Theorem 1 can be adapted to a new context in which $k_{m}$-step TD ($\Delta$) is employed for each $D_{m}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning. Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{m}$) is consistent across all time scales (i.e., $k_{m}=k,\forall m$).
Although these methods could theoretically be equivalent, such equivalence is not common. Generally, for equivalence to be maintained, the learning rate must remain consistent across all time scales. This represents a significant limitation, as shorter time scales, which involve fewer future steps, can be learned more rapidly than longer time scales that require consideration of additional steps ahead. Furthermore, in practical applications, particularly involving nonlinear function approximation such as deep neural networks, adaptive optimizers are frequently employed[30, 28]. These optimizers modify the learning rate in accordance with the complexity of the learning task, which is contingent upon the characteristics of the delta estimator and its target. The effective learning rate varies across different time scales, making it challenging to apply a uniform learning rate in this context.
Besides the learning rate, the method of dissecting the action-value function into distinct components (particularly, $D_{m}$ components) offers benefits that conventional TD learning, including non-delta estimators or elevated $D_{m}$ components, does not possess. The capacity to employ diverse k-step returns (or $\lambda$-return) across many time-scales offers improved control and flexibility in the learning process. If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach may decrease variance while introducing bias, as shorter time-scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9]. In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$. For each iteration n, also called phase n, to get the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a), we average over the trajectories, resulting in:
$\hat{Q}_{\eta,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(x)}+\eta^{k}\hat{Q}_{\eta,n-1}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (27)
Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$
Theorem 3.
Inspired by Kearns & Singh[9], for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. with probability $1-\delta$,
$\Delta_{n}^{\hat{Q}_{\eta}}\leq\epsilon\bigg{(}\underbrace{\frac{1-\eta^{k}}{1-\eta}\bigg{)}}_{\text{variance term}}+\underbrace{\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}}_{\text{bias term}}$ (28)
The proof is presented in appendix PROOFS.
In Eq. 28, A variance term ($\frac{1-\eta^{k}}{1-\eta}$) is arising due to sampling error from rewards collected along trajectories. In particular, $\epsilon$ bounds the deviation of the empirical average of rewards from the true expected reward. Second term ($\eta^{k}\Delta_{n-1}^{\hat{Q}_{\eta}}$) in Eq. 28 is a bias term that arises from bootstrapping off the previous phase’s estimates. Similarly, we consider a phased SARSA version of multi-step $TD(\Delta)$. For each phase n, we update each D as follows:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=1}^{k-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})Q_{\eta_{m-1}}(s_{n+k}^{(x)},a_{n+k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{n+k}^{(x)},a_{n+k}^{(x)})$ (29)
We will now ascertain the highest limit of the phased error. The cumulative errors of each D component, denoted as $TD(\Delta)$, is expressed as $\sum_{m=0}^{M}\Delta_{n}^{m}$, where $\Delta_{n}^{m}=\max_{s,a}\bigg{\{}\bigg{|}\hat{D}_{m}(s,a)-D_{m}(s,a)\bigg{|}\bigg{\}}$. Fundamental differences between SARSA and Q-learning regarding the error bound: In SARSA, the error is ascertained by the agent’s actual actions, indicating that the Q-values are modified based on the state-action pairings encountered by the agent during exploration. The inaccuracy in Q-learning arises from the greedy action, defined as the action that maximizes the Q-value for the subsequent state, based on the premise that the agent consistently behaves optimally.
Theorem 4.
Inspired by Romoff et al.[8], assume that $\eta_{0}\leq\eta_{1}\leq\eta_{2}\leq.....\leq\eta_{m}=\eta$ and $K_{0}\leq K_{1}\leq...\leq K_{m}=K$, for any $10<\delta<$ 1, let $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$, with probability $1-\delta$,
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\underbrace{\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}}_{\text{bias introduction}}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (30)
The proof is presented in appendix PROOFS.
The author examines the comparison of the limits for phased TD($\lambda$) in theorem 3 with those for phased TD($\Delta$) in theorem 4. This comparison demonstrates that phased TD($\Delta$) facilitates variance reduction equivalent to $\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\leq 0$ but introduces a potential bias quantified as $\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}\geq 0$. Utilizing phased TD($\Delta$) to decrease $k_{m}$ synchronizes updates more effectively with recent actions, hence lowering variance from high-discounted return components. Shortened $k_{m}$ values accrue bias more swiftly, as SARSA bootstraps from current policy estimations that may vary over phases. It is observed that when all $k_{m}$ values are identical, both algorithms yield the same upper bound.
In SARSA, the return estimate encompasses subsequent states and actions dictated by the policy. The expected discounted return over T steps closely approximates the infinite-horizon discounted expected return after $\mathrm{T}$, where $\mathrm{T}\approx\frac{1}{1-\eta}$[31]. Consequently, we can effectively simplify $k_{m}$ for any $\eta_{m}$ such that $k_{m}\approx\frac{1}{1-\eta_{m}}$, thereby adhering to this principle. Therefore, with $\mathrm{T}$ samples, similar to TD($\Delta$)[8], SARSA employing TD($\Delta$) utilizes $k_{m}\approx\frac{1}{1-\eta_{m}}$ to establish a suitable balance between bias and variance over all time scales significantly less than $\mathrm{T}$. By establishing $k_{m}$ for each $\eta_{m}$, where $\eta_{m}^{\frac{1}{1-\eta_{m}}}\leq\frac{1}{e}$, the approach guarantees that the effective horizon and variance are maintained within limits without requiring extensive parameter optimization. Each increase in $D_{m}$ facilitates the doubling of the effective horizon, hence ensuring logarithmic scaling with the quantity of action-value functions. SARSA can utilize this to modify additional parameters in alignment with the specified time scales, enhancing the algorithm’s efficiency in calculating long-term rewards without added complexity.
6 Experimental Setup
We experimented with OpenAI Gym control tasks, and Fig.1 shows the relevant environments. The experiments were carried out on a single PC with 16 GB of memory, an Intel Core i7-2600 processor, and a GPU. The system ran 64-bit Ubuntu 22.04.5 LTS. We utilized Python 3.9 and the PyTorch package for development. We also used the OpenAI Gym toolkit, as mentioned by G. Brockman et al. [brockman2016openai]. Ten randomly selected seeds were used in each experiment, and the average outcomes are presented.
(a) FrozenLake-v1
(b) Taxi-v3
(c) Blackjack-v1
(d) CliffWalking-v0
Figure 1: OpenAI Gym control used for all experiments. In order from the left to right: Frozen Lake, Taxi, Blackjack and Cliff Walking.
6.1 Tabular
We utilize Q-value iteration to determine the optimal Q-value function. We employ an open-source version of Value Iteration (VI) derived from the code of Peter Henderson and Wei-Di Chang[value-iteration], guided by the algorithmic principles outlined in Bellman[22] and Sutton & Barto[29]. Although their methodology calculates the optimal value function, we modify it to determine the optimal Q-value function. Upon acquiring the optimal Q-values, we employ Temporal Difference (TD) and SARSA($\Delta$) to derive approximation Q-values via sampling. After 5000 steps, each learning iteration comes to an end. For every setup, we run 200 trials using different random seeds that influence the dynamics of the transition. We implement our algorithms using the code that Kostrikov provided[pytorchrl]. We compare our approach to k-step SARSA, as described in Sutton and Barto’s book, Chapter 7.2[29]. We evaluate SARSA($\Delta$) on multiple RL benchmarks, such as the Ring MDP[9] and deterministic & stochastic environments, to determine its effectiveness. A illustration of the Ring MDP is shown in Fig. 2, which shows a Markov process with five states arranged in a circle. There is a 0.95 chance of moving one state clockwise around the ring and 0.05 chance of staying in the present state at each step. While the other states on the ring offer a reward of 0, two adjacent states offer rewards of +1 and -1, respectively.
Figure 2: Five-state ring MDP as described by Kearns[9]
All experiments utilize the identical parameters established by Romoff et al.[8], which involve supplying a variable number of gamma values, commencing at 0 and incrementing according to $\eta_{m+1}=\frac{\eta_{m}+1}{2}$ until the requisite maximum $\eta_{m}$ is attained. As previously mentioned in Section 5.2, $K_{m}=\frac{1}{(1-\eta_{m})}$ for every m. The baseline consists of a singular estimator with $\eta=\eta_{m}$ and $k=k_{m}$. Fig. 3(a) depicts the efficacy of two methodologies: K-step SARSA ($\Delta$) and K-step SARSA TD, employing a discount factor ($\eta$ = 0.996), k = 64 (a parameter associated with the algorithm’s step count), and diverse values of the learning rate $\alpha$. The y-axis denotes the absolute error, whereas the x-axis signifies the parameter $\alpha$, ranging from 0.0 to 1.0. At $\alpha$=0.0, both methodologies commence at an equivalent error level (about 0.00230). As $\alpha$ increases, K-step SARSA ($\Delta$) initially diminishes its inaccuracy marginally before seeing a consistent rise. K-step SARSA TD, conversely, exhibits a consistent increase in error with minimal fluctuation. Overall course: Relative to K-step SARSA TD, the K-step SARSA ($\Delta$) technique has greater stability and consistently yields superior outcomes (i.e., reduced error) across all values of $\alpha$. Both approaches exhibit sensitivity to variations in $\alpha$, although K-step SARSA TD demonstrates greater sensitivity, evidenced by its more pronounced increase in error.
Fig. 3(b) demonstrates the effectiveness of the two methodologies, K-step SARSA ($\Delta$) and K-step SARSA TD, in relation to the number of k-steps. The y-axis indicates the absolute error, whereas the x-axis represents k-steps, which refers to the number of steps utilized in the algorithms. The range of k-steps extends from 0 to roughly 250. K-step SARSA ($\Delta$) commences with a reduced error at smaller k-steps, which subsequently increases with bigger k-steps. K-step SARSA TD begins with an error level comparable to K-step SARSA ($\Delta$) for low k-steps; however, its error escalates more steeply as k-steps grow. The overall pattern indicates that both techniques exhibit escalating inaccuracy as k-steps rise. K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in all k-step values by exhibiting a reduced error rate. The swift error amplification in K-step SARSA TD indicates that the scheduling method employed in K-step SARSA ($\Delta$) mitigates error accumulation. Furthermore, in this specific MDP, the error escalates with k due to the bias-variance trade-off associated with k-step returns, aligning with the conclusions of Kearns & Singh[9, 8].
Comparison of Figs. 3(a) and 3(b): In the comparison of the two figures, K-step SARSA ($\Delta$) consistently surpasses K-step SARSA TD in both circumstances ($\alpha$ and k-steps). Both approaches exhibit sensitivity to elevated parameter values ($\alpha$ or k-steps). K-step SARSA ($\Delta$) exhibits enhanced stability and attains reduced errors as a result of its scheduling approach. Significantly, errors escalate more rapidly with K-step SARSA TD as k-steps grow in comparison to the increase in $\alpha$.
(a) Error at different learning rates
(b) Error at different k values
Figure 3: Fig.3(a) illustrates the outcomes for $\eta_{M}=0.996$ on the five-state Ring MDP. The error signifies the absolute discrepancy between the estimated and actual discounted Q-value function (precomputed via Q-value iteration), averaged across the complete learning trajectory of 5000 timesteps. Fig.3(b) illustrates the average absolute error for the optimal learning rate at each k-step return, extending beyond the effective planning horizon of $\eta_{M}$.
6.2 OpenAI Gym Environments
The following plots show the results of two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, on the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments from OpenAI Gym. The y-axis shows the average rewards obtained by the algorithms over episodes, and the x-axis shows the number of episodes needed to train the algorithms. Episodes range from 0 to 100,000 episodes. The shaded area surrounding each curve shows the standard error over 10 random seeds. The algorithms both use k= 32, which is the number of steps. The remaining hyperparameter settings are displayed in the table 1.
Table 1: Hyperparameter configurations and their corresponding values.
Hyperparameter Value
seed (10) np.random.randint(40, 45)
Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
k-Step [1, 4, 8, 16, 32, 128, 256]
Number of Episodes 100000
Max Epsilon 0.4
The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig. 4(a) and Fig. 4(b). Fig. 4(a) illustrates the FrozenLake-v1 environment, whereas Fig. 4(b) depicts the Taxi-v3 environment. Fig. 4(a) demonstrates that k-step SARSA ($\Delta$) continually attains superior average rewards relative to k-SARSA TD as training advances. K-step SARSA ($\Delta$) demonstrates considerable enhancement in average rewards over the initial 20,000 episodes, signifying accelerated learning, but k-step SARSA TD exhibits lesser increase in average rewards in comparison to k-step SARSA ($\Delta$) during the same episodes. Following 40,000 episodes, the efficacy of k-step SARSA($\Delta$) stabilizes while consistently surpassing that of k-step SARSA TD, indicating enduring reliability. k-step SARSA regularly exhibits inferior performance relative to k-step SARSA ($\Delta$).
Fig. 4(b) illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance. Fig. 4(b) also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process. The K-Step SARSA TD algorithm encounters difficulties in acquiring an effective policy in this setting compared to K-Step SARSA ($\Delta$). Experiment result show that K-Step SARSA ($\Delta$) significantly outperforms K-Step SARSA TD in the Taxi-v3 environment, achieving much higher average rewards and faster convergence.
(a) FrozenLake-v1
(b) Taxi-v3
Figure 4: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: FrozenLake-v1 and Taxi-v3. The shaded area illustrates the standard error calculated from 10 random seeds.
(a) CliffWalking-v0
(b) Blackjack-v1
Figure 5: Comparative analysis of K-step SARSA ($\Delta$) versus traditional K-step SARSA TD within the OpenAI Gym environments: CliffWalking-v0 and Blackjack-v1. The shaded area illustrates the standard error calculated from 10 random seeds.
The Fig. 5(a) and 5(b) illustrates the learning effectiveness of two variations of the K-Step SARSA algorithm, namely K-Step SARSA ($\Delta$) and K-Step SARSA TD, across 100,000 episodes in the CliffWalking-v0 and Blackjack-v1 environments. Both algorithms are assessed using k=32, a hyperparameter that specifies the number of steps in the updating process. The trend of average rewards in fig. 5(a) demonstrates that both algorithms show an increasing pattern in average rewards as the number of episodes rises, indicating improved policy performance with continued learning. Nevertheless, K-Step SARSA ($\Delta$) often outperforms K-Step SARSA TD, achieving higher average rewards over the episodes.
The red line in fig. 5(a) denoting K-Step SARSA ($\Delta$) demonstrates a considerable enhancement in the initial episodes (about 20,000 episodes), indicating a swifter convergence to optimal policies relative to K-Step SARSA TD. In the 100,000th episode, K-Step SARSA ($\Delta$) attains a reward near -1050, whereas K-Step SARSA TD obtains a slightly lower payout of approximately -1120. This demonstrates the enhanced efficacy of K-Step SARSA ($\Delta$) in this task.
The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in fig. 5(b) that represents K-Step SARSA ($\Delta$). The blue curve in fig. 5(b), which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes. K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards. The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
Finally, when comparing the results of all four environments, the findings demonstrate the constant advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD across all evaluated environments. This superiority is seen through elevated average rewards, accelerated convergence, and enhanced stability in learning outcomes. Environments characterized by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, demonstrate the clear superiority of K-Step SARSA ($\Delta$), but intricate stochastic environments like Blackjack-v1 and Taxi-v3 further highlight its resilience. The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions. In contrast, the performance graph of complex stochastic environments, like Blackjack-v1 and Taxi-v3, appears as a flat line because of the uncertainty and variability in state transitions.
7 Conclusion
With the help of TD($\Delta$), SARSA($\Delta$) improves SARSA by integrating multiple discount factors, leading to better stability and performance. With this approach, we can get a more precise breakdown of action-values while still enjoying the benefits of on-policy learning. In addition to the deterministic and stochastic OpenAI Gym environments, we also conducted trials in a tabular synthetic environment, specifically the five-star ring MDP. In all investigated environments, our experimental results demonstrate that SARSA($\Delta$), our suggested approach, outperforms k-step SARSA TD in terms of greater average rewards, faster convergence, and more stable learning trajectories. The results of this work advance our understanding of multi-step RL algorithms and how they are applied to various sequential decision-making tasks. This information can aid in the creation of robust algorithms for unpredictable, real-world environments in the future. In essence, the Q-Learning version of TD($\Delta$) extends this technique. We decided to take care of it later.
Acknowledgments
Author contributions statement
Data availability
Competing interests
The authors disclose no conflicting interests.
Appendix
PROOFS
Proof: Theorem 1
The proof utilizes induction. Initially, it is necessary to establish that the base case is valid at n=0. This statement is evidently valid given our initial assumption, particularly in the context of zero initialization. At a specific time-step n, we assume the validity of the statement, namely, $\sum_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$. We will now demonstrate that it also holds at the subsequent time step, n+1.
$\sum_{m=0}^{M}\theta_{n+1}^{m}$ $=\sum_{m=0}^{M}\bigg{(}\theta_{n}^{m}+\alpha_{m}\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n})\bigg{)}$
[…]
$=\theta_{n}^{\eta}+\sum_{m=0}^{M}\alpha_{m}\bigg{(}\sum_{k=n}^{\infty}(\lambda_{m}\eta_{m})^{k-n}\delta_{k}^{m}\bigg{)}\phi(s_{n},a_{n})\text{ , Assumption based on induction.}$.
$=\theta_{n}^{\eta}+\alpha\sum_{k=n}^{\infty}(\lambda\eta)^{k-n}\underbrace{\bigg{(}\sum_{m=0}^{M}\delta_{k}^{m}\bigg{)}}_{*}\phi(s_{n},a_{n})\text{ , Because of $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$,$\forall m$}$
To show that $\sum_{m=0}^{M}\theta_{n+1}^{m}=\theta_{n+1}^{\eta}$, we must establish that the term (*) = $\sum_{m=0}^{M}\delta_{k}^{m}$ is equivalent to the conventional TD error $\delta_{k}^{\eta}$.
$\sum_{m=0}^{M}\delta_{k}^{m}$ $=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta_{0}}(s_{k},a_{k})+\sum_{m=1}^{M}\bigg{(}(\eta_{m}-\eta_{m-1})\sum_{q=0}^{m-1}\langle\theta_{n}^{q},\phi(s_{k+1},a_{k+1})\rangle+\eta_{m}\langle\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{n}^{m},\phi(s_{k},a_{k})\rangle\bigg{)}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{m=1}^{M}\bigg{(}\eta_{m}\sum_{q=0}^{m}\theta_{n}^{q}-\eta_{m-1}\sum_{q=0}^{m-1}\theta_{n}^{q}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k},a_{k})\bigg{\rangle}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta_{m}\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\eta_{0}\bigg{\langle}\theta_{n}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=\delta_{k}^{\eta}$ (32)
Proof: Theorem 2
The following is the definition of the Bellman operator:
$\mathrm{T}=r+\eta P$
The discount factor, the expected reward function, and the transition probability operator that is induced by the policy $\pi$ are denoted by the symbols $\eta$, r, and P, respectively. A geometric weighted sum of $\mathrm{T}$ is the definition of the TD($\lambda$) operator, which can be expressed as follows:
$\mathrm{T}=(1-\lambda)\sum_{k=0}^{\infty}\lambda^{k}(\mathrm{T})^{k+1}$ (34)
It may be inferred that $\lambda\in[0,1]$ is necessary for the aforementioned sum to be finite. A corresponding definition of $\mathrm{T}_{\lambda}$ is as follows for every function D:
$\mathrm{T}_{\lambda}D=D+(I-\lambda\eta P)^{-1}(\mathrm{T}D-D)$ (35)
The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $10\leq\lambda\eta<$ 1, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, the Eq. $I-\lambda\eta P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 34 and 35 is lost. This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 35. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 35 can be rewritten as follows:
$\mathrm{T}_{\lambda}$ $=(I-\lambda\eta P)^{-1}(\mathrm{T}D-\lambda\eta PD)$ (36)
The contraction property is demonstrated by examining two action-value functions, $D_{1}$ and $D_{2}$. The distinction between the updates of $\mathrm{T}_{\lambda}D_{1}$ and $\mathrm{T}_{\lambda}D_{2}$ is:
$\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}$ $=(I-\lambda\eta P)^{-1}\bigg{(}\mathrm{T}D_{1}-\mathrm{T}D_{2}-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
$=(I-\lambda\eta P)^{-1}\bigg{(}\eta P(D_{1}-D_{2})-\lambda\eta P(D_{1}-D_{2})\bigg{)}$
$=(I-\lambda\eta P)^{-1}\bigg{(}\eta(1-\lambda)P(D_{1}-D_{2})\bigg{)}$
We derive the following:
$\|\mathrm{T}_{\lambda}D_{1}-\mathrm{T}_{\lambda}D_{2}\|\leq\frac{\eta|1-\lambda|}{1-\lambda\eta}\|D_{1}-D_{2}\|\textit{ where P = 1}$
It is established that $0\leq\lambda\leq 0$ constitutes a contraction. For $\lambda>1$, the following condition must be satisfied:
$\frac{\eta(\lambda-1)}{1-\lambda\eta}<1\Rightarrow\eta\lambda-\eta<1-\lambda\eta$
$\Rightarrow 2\lambda\eta<1+\eta\Rightarrow\lambda<\frac{1+\eta}{2\eta}$ (37)
Consequently, $\mathrm{T}_{\lambda}$ constitutes a contraction if $0\leq\lambda<\frac{1+\eta}{2\eta}$. This directly indicates that for $\eta<1$, $\eta\lambda<1$.
Proof: Theorem 3
[…]
$Q^{\pi}(s,a)=\mathbf{E}\bigg{[}\eta_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q^{\pi}(s_{k},a_{k})\bigg{]}$
Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached. The phased TD(k) update sums the terms $\eta^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\eta^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
[…]
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}$ (38)
If we assume n and the probability of exceeding an $\epsilon$ value is fixed to be atmost $\delta$, we can solve for the resulting value of $\epsilon$:
$2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}=\delta$
$e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}=\frac{\delta}{2}$
Taking the Natural Logarithm
$ln\bigg{(}e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}\bigg{)}}\bigg{)}=ln\bigg{(}\frac{\delta}{2}\bigg{)}$
$\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{n2^{2}}=ln\frac{\delta}{2}$
$\mathit{-}\frac{n\epsilon^{2}}{2}=ln\frac{\delta}{2}$
$\mathit{-}\frac{n\epsilon^{2}}{2}=ln\frac{\delta}{2}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}ln\frac{\delta}{2}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}\bigg{(}ln(\delta)-ln(2)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=\mathit{-}\bigg{(}ln(\delta)+ln(2)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=\bigg{(}ln(2)\mathit{-}ln(\delta)\bigg{)}$
$\frac{n\epsilon^{2}}{2}=ln\frac{2}{\delta}$
$n\epsilon^{2}=2log\frac{2}{\delta}$
$\epsilon^{2}=\frac{2log\frac{2}{\delta}}{n}$
$\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (39)
So by Hoeffding, if we have n samples, with probability at least $1\mathit{-}\delta$,
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2}{\delta}}{n}}$ (40)
Since we are dealing with k different state-action pairs, we use a union bound. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\bigg{)}\leq\epsilon=\sqrt{\frac{2log\frac{2k}{\delta}}{n}}$
We can now assume that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}r_{0}+\eta r_{1}+...+\eta^{k-1}r_{k-1}+\eta^{k}Q_{n}(s_{k},a_{k})\bigg{)}\mathit{-}Q(s,a)$
$=\sum_{i=0}^{k-1}\eta^{i}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (41)
where in the second line we re-expressed the value in terms of a sum of k rewards. We now upper bounded the difference from $E[r_{i}]$ by $\epsilon$ to get
$\hat{Q}_{n+1}(s,a)\mathit{-}Q(s,a)\leq\sum_{i=0}^{k-1}\eta^{l}\epsilon+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$
The variance term arises from the deviation of the empirical average of rewards from the true expected reward:
$\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}$
$\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\bigg{(}\frac{1}{n}\sum_{i=1}^{n}Q_{n}(s_{k}^{i},a_{k}^{i})-\mathbb{E}[Q(s_{k},a_{k})]\bigg{)}$ (42)
The bias term is propagated through bootstrapping:
$\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$
and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
$\Delta^{\hat{Q}_{\eta}}_{n}\leq\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\eta^{k}\Delta^{\hat{Q}_{\eta}}_{n-1}$ (43)
Proof: Theorem 4
Phased $TD(\Delta)$ update rules for $m\geq 1$:
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\bigg{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (44)
We know that according to the multi-step update rule 17 for $m\geq 1$:
$D_{m}(s_{n},a_{n})=\mathbb{E}\bigg{[}\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}r_{i}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}Q_{\eta_{m-1}}(s_{k},a_{k})+\eta_{m}^{k_{m}}D_{m}(s_{k},a_{k})\bigg{]}$ (45)
Then, subtracting the two expressions gives for $m\geq 1$:
$\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$
$+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$ (46)
Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall m\geq 1$, we have:
$\Delta_{n}^{m}\leq\epsilon\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\epsilon\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$ (47)
and $\Delta_{n}^{0}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\eta_{0}^{k_{0}}\Delta_{n-1}^{0}$
Summing the two previous inequalities gives:
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\epsilon\frac{1-\eta_{0}^{k_{0}}}{1-\eta_{0}}+\epsilon\sum_{m=1}^{M}\bigg{(}\frac{1-\eta_{m}^{k_{m}}}{1-\eta_{m}}-\frac{1-\eta_{m-1}^{k_{m}}}{1-\eta_{m-1}}\bigg{)}+\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\underbrace{\epsilon\frac{1-\eta_{M}^{k_{M}}}{1-\eta_{M}}+\epsilon\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta_{m}^{k_{m}}}{1-\eta_{m}}}_{\text{(*)variance term}}+\underbrace{\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}}_{\text{(**)bias term}}$ (48)
Let’s focus now further on the bias term (**)
$\sum_{m=1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\sum_{q=0}^{m-1}\Delta_{n-1}^{q}+\eta_{m}^{k_{m}}\Delta_{n-1}^{m}=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M}\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\Delta_{n-1}^{q}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M}\eta_{m}^{k_{m}}-\sum_{m=q}^{M-1}\eta_{m}^{k_{m+1}}\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\Delta_{n-1}^{q}\bigg{(}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})+(\eta_{M}^{k_{m}}-\eta_{q}^{k_{q+1}})\bigg{)}+\sum_{m=1}^{M}\eta_{m}^{k_{m}}\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\sum_{m=q+1}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}+\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{m}$
$=\sum_{q=0}^{M-1}\sum_{m=q}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$
$=\sum_{m=0}^{M-1}(\eta_{m}^{k_{m}}-\eta_{m}^{k_{m+1}})\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta_{M}^{k_{m}}\sum_{m=0}^{M}\Delta_{n-1}^{m}$ (49)
Finally, we obtain:
$\sum_{m=0}^{M}\Delta_{n}^{m}\leq\underbrace{\epsilon\bigg{(}\frac{1-\eta^{k}}{1-\eta}\bigg{)}+\epsilon\bigg{(}\sum_{m=0}^{M-1}\frac{\eta_{m}^{k_{m+1}}-\eta^{k_{m}}_{m}}{1-\eta_{m}}\bigg{)}}_{\text{variance reduction}}+\underbrace{\sum_{m=0}^{M-1}\bigg{(}\eta^{k_{m}}_{m}-\eta^{k_{m+1}}_{m}\bigg{)}\sum_{q=0}^{m}\Delta_{n-1}^{q}+\eta^{k}\sum_{m=0}^{M}\Delta_{n-1}^{m}}_{\text{bias introduction}}$ (50)